# Interplay of space charge, intrabeam scattering and synchrotron radiation in the Compact Linear Collider damping rings

### v3

@@ VII Impact of vertical dispersion
In rings, the equilibrium vertical emittance is defined by coupling with the horizontal plane and the residual vertical dispersion, which can be induced by quadrupole tilt errors or dipole misalignments and is therefore defined by the vertical emittance tuning of the machine [39]. In the previous sections the presented results did not take into account the vertical dispersion, the impact of which will be addressed here.
Benchmarking simulations for IBS calculations were performed for a lattice with quadrupolar tilt errors, creating vertical dipole magnetic components and thus, vertical dispersion. The applied errors are arbitrary but still consistent with the CLIC requirements, that is an output normalised vertical emittance of below 5 nm. In this case, the original model of BM was used, which takes into account vertical dispersion. Figure 26 shows the results in comparison to the results of the ideal lattice, previously shown in Fig. 4 using the model of Nagaitsev. Vertical dispersion does not contribute significantly in the IBS calculations for the horizontal and longitudinal planes, while for the vertical plane, the output emittance is 34 higher, as by neglecting vertical dispersion, the coupling between the longitudinal and vertical planes is neglected.
To understand the impact of vertical dispersion in the interplay studies during the full cycle of the CLIC DRs, a tracking simulation similar to Section VI.1 was performed taking into account the vertical dispersion in the IBS calculations. Vertical dispersion is produced by arbitrary quadrupolar errors such that the resulting equilibrium parameters from SR and QE are compliant with the target extraction values. Simulation results are presented in Fig. 27, averaged over three runs, where the relative horizontal emittance (blue), vertical emittance (green) and momentum spread (red) are shown, in comparison with the case without vertical dispersion of Fig. 20. Results show that in the presence of SR, the induced vertical dispersion does not affect the horizontal and longitudinal planes, while it produces an extra increase in the vertical steady state emittance. Evidently, the effect of vertical dispersion in the final steady state is insignificant and can be neglected. This allowed the use of Nagaitsev’s IBS model for more efficient computations [62] comparing to the model of BM. In another case where additional errors are present, the impact of the resulting vertical dispersion might be significant and should be studied individually.
@@ VIII Conclusions
A study the incoherent effects of SC, IBS and SR in the context of the CLIC DRs was performed. To do that, two modules for simulating IBS and the SR effects were implemented in PyORBIT in addition to SC. Both modules were successfully benchmarked against analytical calculations and showed excellent agreement.
@@
Simulation studies have been performed to investigate the impact of an excited resonance (here the resonance was chosen) on the achievable beam parameters of the CLIC DRs. It was shown that at the steady state emittances, the beam degradation due to resonance crossing induced by SC alone is relatively weak. This is due to the synchrotron period of the CLIC DRs of about 100 turns which results in resonance side-bands rather than periodic resonance crossing, as demonstrated by the detailed investigation of the SC driven beam dynamics near the resonance. However, particle diffusion and thus emittance growth is strongly enhanced when taking into account also IBS. Though, eventually, the strong damping from SR counteracts this diffusion and the final beam degradation with all three effects combined is of the order of a few percent even for the case of the strong vertical sextupole resonance studied here.
@@
For the first time, combined macro-particle tracking simulations with SC, IBS and SR were performed for the operational scenario of the CLIC DRs, including all parts of the cycle, i.e., starting from injection with large emittances that are damped due to the strong radiation, leading to enhanced SC at the end of the cycle. The nominal working point was used and no errors in the lattice were considered. This is similar to using the full errors of the ring with the corresponding corrections that decrease the error spread back to the systematic errors. The fulfilment of the CLIC target parameters for ultra-high brightness beams is now proven through macro-particle tracking simulations. In the absence of errors, SC does not play any role in the final steady state of the beam, which is then given by the combined effect of SR and IBS alone.
@@
On the other hand, when resonance excitation is enhanced due to errors, it was demonstrated that the choice of the working point is critical. Even for the case of the CLIC DRs with the strong radiation damping and the synchrotron period of about 100 turns, when the tune is set relatively close to an excited resonance. In this case, the beam quality can be degraded as shown for the example of the resonance, for which both the vertical emittance and intensity requirements could not be met. Therefore, the design study of such machines need to include an assessment of resonance excitation in the vicinity of the designed working point and tune footprint evolution along the cycle.
Comparing with a simulation setup with slower synchrotron motion showed that the nominal synchrotron motion of the CLIC DRs makes the beam less sensitive to resonances.
@@
Last, including vertical dispersion in the IBS calculations does not seem to affect the results obtained for the full cycle due to the strong SR.
@@ Appendix A Convergence studies
The third convergence study case follows a different configuration. The initial beam parameters are injection parameters, shown in Table 2, while SC, IBS and SR are included in the simulations. Again, the resonance is strongly excited while the working point is set to . The SC potential is calculated using 200 slices and four different number of macro-particles are chosen; 100, 1000, 5000 and 50000 macro-particles. Figure 30 summarizes the results of the transverse emittances and momentum spread evolution for the last 4000 turns of the cycle that the steady state is reached. The spread is again defined by the standard deviation over three different simulations. In the case of very few macro-particles, e.g., 100 or 1000, the final beam parameters deviate due to the high noise levels in the simulations. Increasing the number of macro-particles to 5000 or above seem to produce accurate results, with less than of relative difference in all three planes when comparing to the case of 50000 macro-particles, while the errors remain small through the cycle. Thus, a particle distribution that consists of 5000 macro-particles, is a good compromise that minimizes the run time of the simulations while maintaining accurate results.

### v4

@@ VII Impact of vertical dispersion
In rings, the equilibrium vertical emittance is dominated by the coupling with the horizontal plane and the residual vertical dispersion, which can be induced by quadrupole tilt errors and dipole misalignments. It is therefore defined through the vertical emittance tuning of the machine [39]. In the previous sections the presented results did not take into account the vertical dispersion, the impact of which will be addressed here.
Benchmarking simulations for IBS calculations were performed for a lattice with quadrupolar tilt errors, creating vertical dipole magnetic components and thus, vertical dispersion. The applied errors are arbitrary but still consistent with the CLIC requirements, that is an output normalised vertical emittance of below 5 nm. In this case, the original model of BM was used, which takes into account vertical dispersion. Figure 26 shows the results in comparison to the ones of the ideal lattice, previously shown in Fig. 4 using the model of Nagaitsev. Vertical dispersion does not contribute significantly in the IBS calculations for the horizontal and longitudinal planes, while for the vertical plane, the output emittance is 34 higher, as by neglecting vertical dispersion, the coupling between the longitudinal and vertical planes is neglected.
To understand the impact of vertical dispersion in the studies during the full cycle of the CLIC DRs, a tracking simulation similar to Section VI.1 was performed taking into account the vertical dispersion in the IBS calculations. In this study, vertical dispersion is produced by arbitrary quadrupolar errors such that the resulting equilibrium parameters from SR and QE are compliant with the target extraction values. Simulation results are presented in Fig. 27, averaged over three runs, where the relative horizontal emittance (blue), vertical emittance (green) and momentum spread (red) are shown, in comparison with the case without vertical dispersion of Fig. 20. Results show that in the presence of SR, the induced vertical dispersion does not affect the horizontal and longitudinal planes, while it produces an extra increase in the vertical steady state emittance. Evidently, the effect of vertical dispersion in the final steady state is insignificant and can be neglected. This allowed the use of Nagaitsev’s IBS model for more efficient computations [62] compared to the model of BM. In cases where additional errors are present, the impact of the resulting vertical dispersion could become significant and should be further studied.
@@ VIII Conclusions
Studies of the incoherent effects of SC, IBS and SR and their interplay were performed in the context of the CLIC DRs. To this end, two modules for simulating IBS and the SR effects were implemented in PyORBIT. Both modules were successfully benchmarked against analytical calculations and showed excellent agreement.
@@
Simulation studies were performed to investigate the impact of an excited resonance (here the resonance was chosen) on the achievable beam parameters of the CLIC DRs. It was shown that at the steady state emittances, the beam degradation due to resonance crossing induced by SC alone is relatively weak. This is due to the synchrotron period of the CLIC DRs of about 100 turns which results in resonance side-bands rather than periodic resonance crossing, as demonstrated by the detailed investigation of the SC driven beam dynamics near the resonance. However, particle diffusion and thus emittance growth is strongly enhanced when IBS is also considered. Eventually, the strong damping from SR counteracts this effect and the beam degradation including SC, IBS and SR effects is of the order of a few percent only even for the case of the strong vertical sextupole resonance studied here.
@@
For the first time, combined macro-particle tracking simulations with SC, IBS and SR were performed for the operational scenario of the CLIC DRs, including all parts of the cycle, i.e., starting from injection with large emittances that are damped due to the strong radiation, leading to enhanced SC at the end of the cycle. The nominal working point was used and no errors in the lattice were considered. This is similar to using the full errors of the ring with the corresponding corrections that decrease the error spread back to the systematic errors. The fulfilment of the CLIC target parameters for ultra-high brightness beams is now demonstrated through macro-particle tracking simulations. In the absence of errors, SC does not play any role in the final steady state of the beam, which is then given by the combined effect of SR and IBS.
@@
On the other hand, when resonance excitation is present due to errors, it was demonstrated that the choice of the working point is critical. Even in the case of the CLIC DRs with the strong radiation damping and the synchrotron period of about 100 turns, significant beam degradation was observed when the tune is set relatively close to the excited resonance. The beam quality degradation was shown for the example of the resonance, where both the vertical emittance and intensity requirements could not be met. Therefore, the design study of such machines need to include an assessment of resonance excitation and the tune spread evolution along the cycle for optimizing the working point. Finally, simulations with slower synchrotron motion showcased how the nominal synchrotron motion of the CLIC DRs makes the beam less sensitive to resonances.
@@
As a last step, vertical dispersion was accounted for in the IBS calculations. It does not significantly affect the results obtained for the full cycle due to the strong SR.
@@ Appendix A Convergence studies
The third convergence study case follows a different configuration. The initial beam parameters are injection parameters, shown in Table 2, while SC, IBS and SR are included in the simulations. Again, the resonance is strongly excited while the working point is set to . The SC potential is calculated using 200 slices and four different number of macro-particles are chosen; 100, 1000, 5000 and 50000 macro-particles. Figure 30 summarizes the results of the transverse emittances and momentum spread evolution for the last 4000 turns of the cycle that the steady state is reached. The spread is again defined by the standard deviation over three different simulations. In the case of very few macro-particles, e.g., 100 or 1000, the final beam parameters deviate due to the high noise levels in the simulations. Increasing the number of macro-particles to 5000 or above produces accurate results, with less than of relative difference in all three planes when comparing to the case of 50000 macro-particles. Thus, a particle distribution that consists of 5000 macro-particles is a good compromise that minimizes the run time of the simulations while maintaining accurate results.